Calculate weighted-average cost of capital for a firm with the following current conditions. Its marginal
tax rate is 30.00%. It has 30 million shares of common stock outstanding that trade for $19.20/share. The
yield to maturity on 10-year US Treasury Bonds is 3.00%. The firm's equity beta is 1.4. The expected return
on the market is 11.00%. The firm's bonds have a 6.50%/yr. coupon rate and $1,000 face value, pay semiannual
coupons, and mature in 15 years. There are 200,000 bonds outstanding and they sell in the open market
for $960.00/bond.
Answer #1

A B C D E F G H Cost 960 Equity Bond MV Weight 2.43% 192000000 25.00% 18.07% 576000000 75.00% 768000000 6.50% Price Years Cou

E F G H I J K MV 960 Cost Equity =D9 Bond =D15 Weight =200000*960 =14/16 =30*10^6*19.2 =15/16 =SUM(14:15) Price Years Coupon